OpenStudy (anonymous):

Please help and show all work, step by step Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

f'(x)= -1 y-2/x+4 = -1 y-2=-x-4 y=-x -2